There were 411 people at a play. Admission was $1.00 for adults and $.75 for children. The receipts were $395.75. How many adults and how many children attended?
Answer by checkley77(12844)

A+C=411 OR A=411-C
1A+.75C=395.75
411-C+.75C=395.75
-.25C=395.75-411
.25C=15.25
C=15.25/.25
C=61 CHILDREN TICKETS WERE SOLD.
A=411-61
A=350 ADULT TICKETS WERE SOLD.
PROOF:
350+.75*61=395.75
350+45.75=395.75
395.75=395.75